OpenStudy (anonymous):

Three friends went shopping. Together, Dan and Jan spent$30. Jan and Van spent $17. Dan and Van spent $25. How much did each friend spend? Please show work.

OpenStudy (allank):

A bit of help: Let the amounts spent by Dan, Jan and Van be d,j and v respectively. d+j=30 j+v=17 d+v=25 Solve that system of equations to get the values of d,j and v.